# CoinFT: A Coin-Sized, Capacitive 6-Axis Force Torque Sensor for Robotic Applications

### v1

@@ I Introduction
Precise force and torque measurement is vital for robots to perform contact-rich tasks safely and effectively. Tasks such as table wiping [1], assembly [2], or palpating soft tissue [3] require the application of force and torque within a specific range—sufficient to complete the task but not so excessive as to cause damage or waste energy. Depending on the application and interaction type, robots performing contact-rich tasks come in various forms, including robotic arms [4], grippers [5], drones [6], and wearable devices [7]. Therefore, equipping these diverse robotic platforms with sensors that can accurately measure force and torque is essential.
@@
Extensive research has been dedicated to developing 6-axis force / torque (F / T) sensors using various transduction methods [8]. Commercially available sensors also exist, such as the Gamma (ATI Industries), and 6-axis F / T sensors from MinebeaMitsumi or ReSense. However, these sensors are challenging to adopt across different robotic platforms, particularly those requiring small, slim form factors and robustness. For example, while commercial sensors can be compact and lightweight, they are often too fragile to reliably withstand impacts from drone crashes or incidental contact on an end-effector. Their thickness also adds unnecessary bulk to haptic devices, making them cumbersome. Additionally, the cost of these sensors can be prohibitive, especially when attempting to sensorize multiple fingers of anthropomorphic robot hands such as the Leap hand [9]. Conversely, existing sensors that are robust and more affordable tend to be bulky and heavy [10], limiting their applicability in compact platforms. For an effective and scalable integration across diverse robotic platforms interacting in unstructured environments, a sensor must be compact, lightweight, robust, and affordable.
@@
In this paper, we introduce CoinFT, a capacitive 6-axis F / T sensor the size of a US quarter dollar, weighing 2 g, with a material cost of less than \$ 10 (Fig. 1 (a)). CoinFT combines all the aforementioned desirable features of a 6-axis F / T sensor—compactness, lightweight, robustness, and affordability—unlocking various contact-rich robot interactions in unstructured environments. One particularly promising application for CoinFT is on drones performing forceful interactions with the environment. Drones have previously executed contact-rich tasks without force feedback [11] in controlled scenarios or through careful human teleoperation [6]. While force sensors have been demonstrated in such tasks, their bulkiness, weight, and fragility make them unsuitable for unstructured, outdoor environments [12]. CoinFT’s design addresses these challenges, making it a viable solution for drones operating in environments where large impulses from contact or crashes are common. Its lightweight nature minimally impacts drone operation time, and its compact size facilitates easy integration into even small drones. In this work, we demonstrate the use of CoinFT on a drone for contact modulation through PID force control and the deployment of objects onto environmental surfaces by applying a controlled range of force.
@@
  • Open sourcing the design, fabrication, and firmware of CoinFT.
@@ II-A Multi-axial Force Torque Sensors
Given its critical role in various contact-rich tasks, significant research has focused on developing multi-axial F / T sensors using diverse transduction methods, including piezoresistive, capacitive, optical, magnetic, and pneumatic techniques [8]. Among these, piezoresistive approaches have been the most extensively explored, typically involving the attachment of multiple strain gauges to precisely machined structures, commonly designed as cross-beam or parallel structures [8]. Piezoresistive sensing offers precise measurement capabilities with a high signal-to-noise ratio (SNR), facilitated by strain gauges and circuitry optimized for SNR amplification [13]. The availability of compact electronics and advanced machining techniques has also enabled the fabrication of compact sensors, which is why commercially available 6-axis F / T sensors predominantly utilize piezoresistive technology (MinebeaMitsumi and ReSense). However, despite these advantages, these sensors are often fragile and susceptible to impact, which can result in plastic deformation or even failure of the sensing medium [14].
@@
Capacitive sensing is another widely adopted approach for designing 6-axis F / T sensors, valued for its potential to achieve a compact form factor while maintaining high sensitivity [8]. Compared to piezoresistive sensors, capacitive sensors can be relatively simpler to fabricate. Multiaxial F / T sensing via capacitive methods is typically achieved in one of two ways: using a single layer of sensing electrodes integrated with specially designed structures that deform and create distinct signal patterns upon input [15, 16], or employing multi-layer structures with electrodes, where each layer is more sensitive to specific subsets of forces and torques, generating unique signal patterns accordingly [17].
@@
Optical sensing methods have been employed in the design of multi-axis F / T sensors by controlling the scattering or reflection of light within the system [18]. These methods can be broadly categorized into light intensity sensing [19], employing Fiber Bragg Grating (FBG) optical fibers [20], and vision-based [21] techniques. In traditional optical methods, variations in reflected light intensity at multiple locations are converted into measurements of force and torque [19]. FBG-based sensing leverages FBG fibers embedded in areas of higher strain within the sensor body, functioning similarly to strain gauges to precisely measure deformation [20]. Vision-based tactile sensing, on the other hand, involves using a compact camera to observe the deformation of a silicone rubber medium, often a robot fingertip, upon contact [21]. A recently popularized approach for force and torque sensing involves monitoring the deformation of specially patterned silicone rubber fingertips and employing data-driven models to infer the applied force and torque [21, 22]. Another method involves tracking the movement of a visual cue with a camera to deduce the load [23, 24]. Alternatively, gel deformation patterns can be remotely observed through an optical fiber, enabling the decoupling of the silicone gel from the sensing camera [25].
@@
Other, less explored approaches in the design of multi-axial F / T sensors include piezoelectric [26], magnetic [27, 28], and pneumatic methods [10]. While piezoelectric effects have been utilized in these sensors, their susceptibility to charge leakage makes them less ideal for measuring static loads. Magnetic sensing techniques, which involve strategically positioning magnets and Hall effect sensors to detect forces and torques across six axes, have also been explored [27]. However, magnetic sensing is inherently vulnerable to external noise sources, such as the Earth’s magnetic field or ferromagnetic objects. Additionally, similar design principles have been applied in pneumatic sensors, where pressure differential patterns are created using an array of barometers [10].
@@ II-B Contact Based Interaction of Drones
Significant research has focused on enabling drones to avoid contact [29], and more recently, efforts have shifted toward allowing drones to interact with their environment through controlled and sustained contact, such as in perching [30] or performing tasks that require contact modulation [31]. Perching, in particular, has garnered attention for its ability to allow drones to remain in an environment without expending energy to stay aloft, thereby significantly increasing their operational time. To facilitate perching on horizontal [32] and vertical surfaces [33], bioinspired grippers [34] and control strategies [11] have been developed. While these approaches show promise for perching, they lack contact sensing, which limits the ability to assess perch stability or react to failures in real-time.
A number of research has explored the use of force-controlled contact with the environment through the integration of load cells [35]. These studies have investigated various applications, such as maintaining a consistent contact force while changing position using hybrid position-force controllers [36], analyzing surface textures by combining load cells with vision-based tactile sensors [12], and attaching objects to surfaces with custom-designed mechanisms [35]. However, the load cells used in these studies, typically commercial 6-axis F / T sensors, present challenges for real-world integration due to their high cost, fragility, and bulky readout electronics. In contrast, CoinFT, with its lightweight, compact design, robustness, and affordability, is ideally suited for drone applications.
@@ III-A CoinFT Design
CoinFT is a capacitive 6-axis force / torque sensor. Its circular sensing area is 20 mm in diameter and approximately 2 mm in thickness (Fig. 1 (c)), having a comparable form-factor to a quarter-dollar coin in the United States currency (Fig. 1 (a)). With a total weight of 2 g, it consists of two 0.8 mm thick rigid PCBs (Printed Circuit Boards) connected by an array of 127 um thick cylindrical pillars cast from silicone rubber (TAP® Silicone RTV Mold-Making System). The radial density of pillars increases towards the edge of the PCB for improved strength against delamination. We employed a dense array of microns-wide pillars despite the added challenges in fabrication to retain sensitivity. Wider structures will increase effective stiffness of the pillar layer, which will be discussed in detail in Section III-D.
@@
Figure 2: CoinFT working principle. The microcontroller firmware is programmed to switch the pair of rigid PCBs between two different configurations - normal mode and shear mode. The capacitance signals collected from these two configurations produce a unique pattern under different force and torque inputs.
Each rigid PCB is multilayered, with four quadrants of a pair of comb-shaped electrodes on one side and a plane of electrode for passive shielding on the other side (Fig. 1 (b), (c)). The upper sensing layer has eight physical electrodes while the lower sensing layer houses two physical electrodes (Fig. 1 (b)). It is by leveraging this design with the unique capabilities of the microcontroller (Cypress, PSoCTM 4100S) on CoinFT, that a pair of PCBs become sensitive to 6-axis F / T, which will be discussed in detail in Section III-B. The top shield layer made of a 127 um thick fPCB (flexible Printed Circuit Board) is then added to provide additional passive shielding, as the via holes in the upper sensing layer can introduce noise from stray capacitance in the environment (Fig. 1 (b), (c)). Each PCB layer has structures for alignment on its sides, which is later removed during the final stages of fabrication.
@@ III-B Working Principle
CoinFT measures 6-axis force and torque by observing the changes in capacitance signal patterns between the upper sensing layer and the lower sensing layer (Fig. 1). Capacitance is expressed as equation (1)
@@
In the normal mode, the pair of electrodes in each quadrant in the upper sensing layer are internally connected to function as a single electrode, effectively reducing the eight physical electrodes to four: Z1, Z2, Z3, and Z4 (Fig. 2). The two electrodes in the lower sensing layer are reconfigured to connect to ground. This design allows the normal mode to be sensitive to inputs that changes $d$, such as Fz, Mx, and My, but not to changes in $A$, which results from Fx, Fy, and Mz. Under Fz, all normal mode electrodes (Z1, Z2, Z3, and Z4) show a rise in signal, while under Mx or My, a differential pattern between pairs of normal mode electrodes is created (Fig. 2).
In the shear mode, each electrode in the upper sensing layer are kept as independent sensing electrodes ($X1$, $X2$, $X3$, $X4$, $Y1$, $Y2$, $Y3$, and $Y4$), while the electrodes in the lower sensing layer are configured such that the ground and shield electrodes form an alternating pattern (Fig. 2). The shield electrode actively mitigates interference from stray capacitance by being excited to the same electrical potential as the sensing electrodes, an innate capability of PSoCTM 4100S. This design allows the shear mode to be sensitive primarily to shear inputs such as $Fx$, $Fy$, and $Tz$. Under a shear pressure, the relative position of the upper sensing layer and the lower layer shifts laterally, changing $A$ between a sensing electrode and ground electrode. Due to the negligible capacitance formation between a sensing electrode and a shield electrode, the relative position change results in a differential signal pattern in the shear mode. The comb shaped electrodes are populated in a way that allows the first and third quadrant of the shear mode is sensitive to loads in the $X$ direction, while the second and fourth quadrants are sensitive to the $Y$ direction. Under torsion ($Tz$), all four quadrants in the shear mode show a differential signal pattern. Unlike the normal mode configuration where there is negligible cross-modal sensitivity by design, there is cross-coupling in the shear mode configuration. This effect and its implications will be discussed in detail Section III-D1. The unique signal pattern generated from the normal and shear mode configurations upon loads in different axes makes CoinFT, a pair of rigid PCBs, a 6-axis F / T sensor.
@@ III-C Sensor Fabrication
Figure 3: The fabrication process of CoinFT. (a) Fresh uncured silicone is spread on a UV laser cut mask that is placed on an acrylic plate with alignment pins. (b) A primed lower sensing layer PCB and an acrylic plate for equal pressure distribution is stacked with a 3.1 kg weight. (c) With a 0.1 kg weight, the assembly is cured inside a pressurized chamber. (d),(e) Once the mask and acrylic plates are removed, the pillar layer is complete. (f) On a primed upper sensing layer PCB, fresh uncured silicone is spread. (g) The silicone layer is made thin and uniform through spin-coating. (h) The lower sensing layer with pillars is assembled with the upper sensing layer through precise distance control by adding spacers in between. A 0.1 kg weight and an acrylic plate for pressure distribution is added. (i) The top shield layer is attached using an adhesive. (j) The cross section of the pillar layer shows desirable bonding of the two PCB layers.
@@ III-D Sensor Characterization
A quantitative analysis using finite element analysis (FEA) and real sample testing has been conducted to understand the behavior of CoinFT. The FEA involved a structural analysis using ANSYS (Fig. 4 (h)) to investigate the deformation of the pillar array under different loading conditions, and an electrostatics analysis in COMSOL (Fig. 4 (i)), to study the change in capacitance in the sensing electrodes resulting from pillar deformation. The pillar mechanics was modelled using a linear elastic model with a high deformation option to account for the geometric nonlinearities. The Gent model [37] was applied to convert the shore hardness of the silicone rubber into Young’s modulus. With the assumption that the silicone rubber is incompressible, the effective Young’s modulus of each pillar was derived using a disk model of an elastomer bonded between two plates, as modeled by Haddow et al. [38], shown in (2).
@@
$E_{e}$ is the effective Young’s modulus, $E$ is the Young’s modulus of the material, and $\eta$ is the ratio between the height and radius of the pillars.
The dielectric constant of the silicone rubber was assumed to be 3.0, a common value for silicone rubber [39], while the dielectric constant of air was taken as 1.0. In order to verify the reliability of our FEA model, we compared the real and simulated response of CoinFT to normal force (Fig. 4 (d)), where the load for the real sample was given up to 10 N at 0.1 Hz. The FEA model closely matched the actual sample response, accurately capturing the stiffening effects due to pillar deformation at higher loads, and effectively representing the true behavior of CoinFT (Fig. 4 (d), (g)).

### v2

@@ I Introduction
Precise force and torque measurement is vital for robots to perform contact-rich tasks safely and effectively. Tasks such as table wiping [1], assembly [2], or palpating soft tissue [3] require the application of force and torque within a specific range—sufficient to complete the task but not excessive, as to cause damage or waste energy. Depending on the application and interaction type, robots performing contact-rich tasks come in various forms, including robotic arms [4], grippers [5], drones [6], and wearable devices [7]. Therefore, it is highly desirable to develop a compact and low cost solution that can serve a wide variety of applications.
@@
Extensive research has been dedicated to developing 6-axis force / torque (F / T) sensors using various transduction methods [8]. Commercially available sensors also exist, such as the Gamma™(ATI Industries), and 6-axis F / T sensors from MinebeaMitsumi™or ReSense™. However, these sensors are challenging to adopt across different robotic platforms, particularly those requiring both robustness and small, slim form factors. For example, while commercial sensors can be compact and lightweight, they are often too fragile to reliably withstand impacts from drone crashes or incidental contact on an end-effector. Their thickness also adds unnecessary bulk to haptic devices, making them cumbersome. Additionally, the cost of these sensors can be prohibitive, especially when attempting to sensorize multiple fingers of anthropomorphic robot hands such as the Leap hand [9]. Other sensors that are robust and more affordable tend to be bulky [10], limiting their applicability in compact platforms.
In this paper, we introduce CoinFT, a capacitive 6-axis F / T sensor the size of a US quarter dollar, weighing 2 g, with a material cost of less than \$ 10 (Fig. 1 (a)). CoinFT combines the aforementioned desirable features, unlocking various contact-rich robot interactions in unstructured environments.
@@
One particularly promising application for CoinFT is on drones performing forceful interactions with the environment. Drones have previously executed contact-rich tasks without force feedback [11] in controlled scenarios or through careful human teleoperation [6]. While force sensors have been demonstrated in such tasks, their size, weight, and fragility make them unsuitable for unstructured, outdoor environments [12]. CoinFT’s design addresses these challenges, making it a viable solution for drones operating in environments where large impulses from contact or crashes are common. Its lightweight nature minimally impacts drone operation time, and its compact size facilitates easy integration into even small drones. In this work, we demonstrate the use of CoinFT on a drone for contact modulation through PID force control and the deployment of objects onto environmental surfaces by applying a controlled range of force.
@@
  • Open-sourcing the design, fabrication, and firmware of CoinFT.
@@ II-A Multi-axial Force Torque Sensors
Given the role of forces in contact-rich tasks, significant research has focused on developing multi-axial F / T sensors using diverse transduction methods, including piezoresistive, capacitive, optical, magnetic, and pneumatic techniques [8]. Among these, piezoresistive approaches have been the most extensively explored, typically involving the attachment of multiple strain gauges to precisely machined structures, commonly designed as cross-beam or parallel structures [8]. Piezoresistive sensing offers precise measurement capabilities with a high signal-to-noise ratio (SNR), facilitated by strain gauges and circuitry optimized for SNR amplification [13]. The availability of compact electronics and advanced machining techniques has also enabled the fabrication of compact sensors, which is why commercially available 6-axis F / T sensors predominantly utilize piezoresistive technology (e.g. MinebeaMitsumi™and ReSense™). However, these sensors are often fragile and susceptible to impact, which can result in plastic deformation or even failure of the sensing medium [14].
@@
Capacitive sensing is another widely adopted approach for designing 6-axis F / T sensors, valued for its potential to achieve a compact form factor while maintaining high sensitivity [8]. Compared to piezoresistive sensors, capacitive sensors can be relatively simpler to fabricate. Multiaxial F / T sensing via capacitive methods is typically achieved in one of two ways: using a single layer of sensing electrodes integrated with specially designed structures that deform and create distinct signal patterns upon input [15, 16, 17], or employing multi-layer structures with electrodes, where each layer is sensitive to specific subsets of forces and torques, generating unique signal patterns accordingly [18].
@@
Optical sensing methods have been employed in the design of multi-axis F / T sensors by controlling the scattering or reflection of light within the system [19]. These methods can be broadly categorized into light intensity sensing [20], employing Fiber Bragg Grating (FBG) optical fibers [21], and vision-based techniques. In traditional optical methods, variations in reflected light intensity at multiple locations are converted into measurements of force and torque [20]. FBG-based sensing leverages FBG fibers embedded in areas of higher strain within the sensor body, functioning similarly to strain gauges to precisely measure deformation [21]. Vision-based tactile sensing involves using a compact camera to observe the deformation of a silicone rubber medium, often a robot fingertip, upon contact [22, 23]. A recently popularized approach for force and torque sensing involves monitoring the deformation of specially patterned silicone rubber fingertips and employing data-driven models to infer the applied force and torque [23, 24, 25]. Another method involves tracking the movement of a visual cue with a camera to deduce the load [26, 27]. Alternatively, gel deformation patterns can be remotely observed through an optical fiber, enabling the decoupling of the silicone gel from the sensing camera [28].
@@
Other, less explored approaches in the design of multi-axial F / T sensors include piezoelectric [29], magnetic [30, 31], and pneumatic methods [10]. While piezoelectric effects have been utilized in these sensors, their susceptibility to charge leakage makes them less ideal for measuring static loads. Magnetic sensing techniques, which involve strategically positioning magnets and Hall effect sensors to detect forces and torques across six axes, have also been explored [30]. However, magnetic sensing is inherently vulnerable to external noise sources, such as the Earth’s magnetic field or nearby ferromagnetic objects. Additionally, similar design principles have been applied in pressure sensors, where differential patterns are created using an array of barometers [10].
@@ II-B Contact Based Interaction of Drones
Significant research has focused on enabling drones to avoid contact [32], and more recently, efforts have shifted toward allowing drones to interact with their environment through controlled and sustained contact, such as in perching [33] or performing tasks that require contact modulation [34]. Perching, in particular, has garnered attention for its ability to allow drones to remain in an environment without expending energy to stay aloft, thereby significantly increasing their operational time. To facilitate perching on horizontal [35] and vertical surfaces [36], bioinspired grippers [37] and control strategies [11] have been developed. While these approaches show promise for perching, they typically lack contact sensing, which limits the ability to assess perch stability or react to failures in real-time.
Recent research has explored the use of force-controlled contact with the environment through the integration of load cells [38]. These studies have investigated various applications, such as maintaining a consistent contact force while changing position using hybrid position-force controllers [39], analyzing surface textures by combining load cells with vision-based tactile sensors [12], and attaching objects to surfaces with custom-designed mechanisms [38]. However, the load cells used in these studies, typically commercial 6-axis F / T sensors, present challenges for real-world integration due to their high cost, fragility, and bulky readout electronics. In contrast, CoinFT, with its lightweight, compact design, robustness, and affordability, is ideally suited for drone applications.
@@ III-A CoinFT Design
CoinFT is a capacitive 6-axis force / torque sensor. Its circular sensing area is 20 mm in diameter and approximately 2 mm in thickness (Fig. 1 (c)), having a comparable form-factor to a quarter-dollar coin in the United States currency (Fig. 1 (a)). With a total weight of 2 g, it consists of two 0.8 mm thick rigid PCBs (Printed Circuit Boards) connected by an array of 127 um tall cylindrical pillars cast from silicone rubber (TAP® Silicone RTV Mold-Making System). The radial density of pillars increases towards the edge of the PCB for improved strength against delamination. We employ a dense array of microns-wide pillars, despite the added challenges in fabrication, to retain sensitivity. Wider structures will increase effective stiffness of the pillar layer, which will be discussed in detail in Section III-D.
@@
Figure 2: CoinFT working principle. The microcontroller firmware is programmed to switch the pair of rigid PCBs between two different configurations - normal mode and shear mode. The capacitance signals collected from these two configurations produce unique patterns under different force and torque inputs.
Each rigid PCB is multilayered, with four quadrants of a pair of comb-shaped electrodes on one side and a plane electrode for passive shielding on the other side (Fig. 1 (b), (c)). The upper sensing layer has eight physical electrodes while the lower sensing layer houses two physical electrodes (Fig. 1 (b)). It is by leveraging this design with the unique capabilities of the microcontroller (Cypress, PSoCTM 4100S) on CoinFT, that a pair of PCBs becomes sensitive to six axes of force and torque, which will be discussed in detail in Section III-B. The top shield layer made of a 127 um thick fPCB (flexible Printed Circuit Board) is then added to provide additional passive shielding, as the via holes in the upper sensing layer can introduce noise from stray capacitance in the environment (Fig. 1 (b), (c)). Each PCB layer has structures for alignment on its sides, which are later removed during the final stages of fabrication.
@@ III-B Working Principle
CoinFT measures 6-axis force and torque by observing the changes in capacitance signal patterns between the upper sensing layer and the lower sensing layer (Fig. 1). Capacitance is expressed as:
@@
In the normal mode, the pair of electrodes in each quadrant in the upper sensing layer are internally connected to function as a single electrode, effectively reducing the eight physical electrodes to four: Z1, Z2, Z3, and Z4 (Fig. 2). The two electrodes in the lower sensing layer are reconfigured to connect to ground. Referring to (1), this design allows the normal mode to be sensitive to inputs that changes $d$, such as Fz, Mx, and My, but not to changes in $A$, which results from Fx, Fy, and Mz. Under Fz, all normal mode electrodes (Z1, Z2, Z3, and Z4) show a rise in signal, while under Mx or My, a differential pattern between pairs of normal mode electrodes is created (Fig. 2).
In the shear mode, electrodes in the upper sensing layer are kept as independent sensing electrodes ($X1$, $X2$, $X3$, $X4$, $Y1$, $Y2$, $Y3$, and $Y4$), while the electrodes in the lower sensing layer are configured such that the ground and shield electrodes form an alternating pattern (Fig. 2). The shield electrode actively mitigates interference from stray capacitance by being excited to the same electrical potential as the sensing electrodes, an innate capability of PSoCTM 4100S. This design allows the shear mode to be sensitive primarily to shear inputs such as $Fx$, $Fy$, and $Tz$. Under a shear pressure, the relative position of the upper sensing layer and the lower layer shifts laterally, changing $A$ between a sensing electrode and ground electrode. Due to the negligible capacitance formation between a sensing electrode and a shield electrode, the relative position change results in a differential signal pattern in the shear mode. The comb-shaped electrodes are populated in a way that allows the first and third quadrant of the shear mode to be sensitive to loads in the $X$ direction, while the second and fourth quadrants are sensitive to the $Y$ direction. Under torsion ($Tz$), all four quadrants in the shear mode show a differential signal pattern. Unlike the normal mode configuration where there is negligible cross-modal sensitivity by design, there is cross-coupling in the shear mode configuration. This effect and its implications will be discussed in detail Section III-D1. The unique signal pattern generated from the normal and shear mode configurations upon loads in different axes makes CoinFT a 6-axis F / T sensor.
@@ III-C Sensor Fabrication
Figure 3: The fabrication process of CoinFT. (a) Fresh uncured silicone is spread on a UV laser cut mask that is placed on an acrylic plate with alignment pins. (b) A primed lower sensing layer PCB and an acrylic plate for equal pressure distribution is stacked with a 3.1 kg weight. (c) With a 0.1 kg weight, the assembly is cured inside a pressurized chamber. (d),(e) Once the mask and acrylic plates are removed, the pillar layer is complete. (f) On a primed upper sensing layer PCB, fresh uncured silicone is spread. (g) The silicone layer is made thin and uniform through spin-coating. (h) The lower sensing layer with pillars is assembled with the upper sensing layer through precise distance control by adding spacers in between. A 0.1 kg weight and an acrylic plate for pressure distribution are added. (i) The top shield layer is attached using an adhesive. (j) A horizontal cross section of the pillar layer shows desirable bonding of the two PCB layers.
@@ III-D Sensor Characterization
A quantitative analysis using finite element analysis (FEA) and real sample testing has been conducted to understand the behavior of CoinFT. The FEA involved a structural analysis using ANSYS (Fig. 4 (h)) to investigate the deformation of the pillar array under different loading conditions, and an electrostatics analysis in COMSOL (Fig. 4 (i)), to study the change in capacitance in the sensing electrodes resulting from pillar deformation. The pillar mechanics was modelled using a linear elastic model with a high deformation option to account for the geometric nonlinearities. The Gent model [40] was applied to convert the shore hardness of the silicone rubber into Young’s modulus. With the assumption that the silicone rubber is incompressible, the effective Young’s modulus of each pillar was derived using a disk model of an elastomer bonded between two plates, as modeled by Haddow and Ogden [41], shown in (2):
@@
where $E_{e}$ is the effective Young’s modulus, $E$ is the Young’s modulus of the material, and $\eta$ is the ratio between the height and radius of the pillars.
The dielectric constant of the silicone rubber was assumed to be 3.0, a common value for silicone rubber [42], while the dielectric constant of air was taken as 1.0. In order to verify the reliability of our FEA model, we compared the real and simulated response of CoinFT to normal force (Fig. 4 (d)), where the load for the real sample was given up to 10 N at 0.1 Hz. The FEA model closely matched the actual sample response, accurately capturing the stiffening effects due to pillar deformation at higher loads, and effectively representing the behavior of CoinFT (Fig. 4 (d), (g)).